As a technician in a large pharmaceutical research firm, you need to produce 200. mL of a potassium dihydrogen phosphate buffer solution of pH = 7.07. The pKa of H2PO4− is 7.21. You have the following supplies: 2.00 L of 1.00 M KH2PO4 stock solution, 1.50 L of 1.00 M K2HPO4 stock solution, and a carboy of pure distilled H2O. How much 1.00 M KH2PO4 will you need to make this solution? (Assume additive volumes.)
Answer #1

There is a typo error, we are provided with

2.00L of 1.00M KH2PO4 olution and 1.50L of 1.00M K2HPO4 solution.

Using H. H equation,

pH = pKa + log [anion] / [acid]
7.07= 7.21 + log [anion] / [acid]
-0.14=log [anion] / [acid]
or, 0.724435= [anion] / [acid]

Ratio of concentrations in a solution, (200 ml), is the same as the ratio of moles in that volume.

arbitrarially choosing such that one to be added "X" ml of it , & the other becomes that its been added "250-X" ml

0.724436= (250 ml - X ml)(1.00 M K2HPO4) / (X ml)(1.00 M KH2PO4)

0.724436 = (250 - X) / X

After solving X = 144.97 ml

Thus we mix 144.97 ml of 1.00 MKH2PO4 with 105.03 ml of 1.00 M K2HPO4 to make 250 ml Solution